# A Unified Approach for Multi-granularity Search over Spatial Datasets

### v2

@@ Abstract
Index Terms:
Spatial dataset, Dataset search, Data point search, Multi-granularity search, Distance metrics.
I Introduction
The development of sensor technology, GPS-enabled mobile devices, and wireless communication brings prosperity to spatial data, and the emergence of a large number of spatial data has fueled the demand for searching spatial data [28, 16, 45, 78, 62, 70, 27]. Since a spatial dataset usually consists of a set of spatial data points [70], various spatial dataset search [70, 41, 42] and data point search systems [9, 22, 29] have been developed. For example, Foursquare [9] helps users find the best location points in any city around the world. OpenStreetMap [8] provides search functionality over a vast collection of spatial datasets contributed by users worldwide. However, they only support keyword-based search over published dataset metadata [19, 82, 23].
The published spatial metadata is often missing or incomplete [1], rendering keyword-based queries insufficient for helping users locate datasets or data points. Therefore, various query operations that rely on raw data have been developed to identify spatial datasets and data points. For example, existing spatial dataset search engines treat a dataset as the smallest return unit. In contrast, in the existing spatial data point search engines, a data point is the smallest return unit. In what follows, we use a traffic analysis example to introduce these two types of searches.
Figure 1: The illustration of our multi-granularity search research over spatial datasets, which seamlessly integrates the dataset search (including (1) and (2) based on three distance metrics) and data point search (including (3) and (4)).
@@ I Introduction
Consider a scenario where traffic analysts want to analyze traffic patterns in a metropolitan area. Firstly, they can define a query range of $R$ [1, 78, 17] to locate all relevant traffic datasets within that area for analysis, as illustrated on the left side of Figure1(a). Then, to identify the most efficient commuting routes, the analysts can input a specific traffic dataset $Q$ and search for datasets similar to $Q$ based on defined similarity measures [70, 65, 66] for data analysis, illustrated on the right side of Figure 1(a). This search manner is known as exemplar search [55, 73, 64, 41, 44, 71]. Furthermore, if the analysts want to find points of interest, such as gas stations or rest areas along the searched routes $D$, they can specify a query range $R$ along with the relevant traffic dataset to retrieve these points, as depicted on the left side of Fig.1(b)[9, 62, 79, 53]. Additionally, on the right side of Figure1(b), the analysts can identify the nearest gas stations for each location in the dataset $Q$ [53, 79, 29, 25, 26], which helps commuters find convenient stops along their routes [74, 72].
@@
What is the Problem? Despite the efforts made in the field of spatial data search systems, existing systems focus mainly on one type of search. There is currently no system that has successfully integrated all types of searches into a unified search framework. In addition, the existing systems, such as Auctus [1] and OpenStreetMap [8], are relatively simple in spatial data search and do not take into account the similarity measurement of spatial location distribution characteristics. Our study, depicted in Figure 1, considers all searches at both the coarse-grained spatial dataset level and the more fine-grained data point level. In spatial dataset search, we integrate range-based dataset search and top-$k$ exemplar dataset search based on several distance metrics.
@@
Figure 2: The illustration of three distance metrics, where (a) shows that the IA of $Q$ and $D_{1}$ is the size of their minimum bounding rectangles’ overlapping areas, i.e., the rectangle filled with red lines, (b) shows that the GBO of $Q$ and $D_{1}$ is 1 since there is one overlapping cell between $Q$ and $D_{1}$, and (c) shows that the Haus of $Q$ and $D_{1}$ is the maximum nearest neighbor distance, i.e., the distance from the point $p_{1}$ of $Q$ to the point $p_{1}^{\prime}$ of $D_{1}$.
@@
As shown in Figure 2, the intersecting area (IA[39, 28] is useful when assessing the overlap between public transportation routes, which can help identify congestion points to optimize traffic management. Compared to IA, grid-based overlap (GBO[45] can capture the density of traffic in the urban grid in a finer granularity and optimize traffic routes. In addition, calculating the Hausdorff distance (Haus[59, 46, 30] can help analysts understand the accessibility of bus stops for residents (the formal definition is in Section III-B). Moreover, building upon the outcomes of the dataset search, we further enhance our capabilities to facilitate a finer-grained data point search process, wherein the results obtained from the dataset search are fed to the subsequent data point search.
@@
What are the Challenges? Our goal is to design a multi-granularity spatial data search system that seamlessly supports coarse-grained dataset search based on multiple distance metrics and fine-grained data point search. There are several challenges: Firstly, A naïve way is to build separate indexes for each query type. However, this method incurs significant costs in both storage and index construction time. Furthermore, the presence of outliers in spatial datasets, such as erroneous data points caused by GPS errors, can severely affect the accuracy and efficiency of similarity computation [63]. Finally, although the index can quickly locate data matching query criteria, it does not effectively reduce the number of candidate points to examine. Thus, there is an urgent need for efficient acceleration strategies to speed up a rich set of searches.
@@
How the Challenges are Addressed? To address the challenges above, we develop a multi-granularity spatial data search system, called Spadas. Specifically, we design a unified index [67, 37] that can effectively organize both datasets and data points, enabling efficient pruning for various search scenarios. In addition, to further enhance system performance, we propose a set of novel and highly efficient pruning mechanisms, such as efficient bound estimation or error-bounded approximation techniques, ensuring that irrelevant candidates are filtered early in the process, thereby reducing computational overhead and improving overall system performance. Our main contributions are summarized as follows.
@@
  • We propose a multi-granularity spatial data search system, called Spadas, to efficiently support coarse-grained dataset search based on multiple distance metrics and fine-grained data point search (see Section IV). We also implement an online Spadas system [^1] http://swangdbs.com/prototype/spadas/ that can be accessed by users.
@@
  • To support both searches with different query types and distance metrics, we specifically design a unified index to enhance the efficiency of data search operations by efficiently organizing spatial data, while mitigating the influence of outliers on the accuracy of similarity measures between datasets (see Section V).
@@
  • To accelerate multiple types of data search operations, we propose a range of pruning mechanisms over the unified index, including tighter lower and upper bounds, approximation technique with error bound, and batch pruning, that effectively filter candidates and thus significantly reduce computational overhead (see Section VI).
@@
  • We conduct extensive experiments on six real-world data repositories, and the results show that our proposed Spadas can achieve orders of magnitude speedup than the state-of-the-art algorithms in dataset search and data point search. The case studies validate the effectiveness of our search system (see Section VII).
II Related Work
@@
II-A Spatial Dataset Search
@@
Overlap-based Similarity Search.  For a user with a query dataset, the overlap-based similarity search ranks candidate datasets based on the size of their overlap area. Computing the overlapping area between two Minimum Bounding Rectangles (MBRs) is a commonly used and straightforward measure. Frontiera et al. [30] proposed several variants by considering the query and the dataset’s range. Vasconcelos et al. [27] and Degbelo et al. [28] applied it as the main ranking metric for an open government dataset search portal. More recently, with the rise of grid partition technology [78, 68], a more fine-grained similarity measure based on the number of overlapping grid cells began to appear, i.e., modeling each dataset as a grid with points inside by dividing the space equally and using the intersection of the grids as the similarity.
@@
Point Set Similarity Search.  Point set similarity search mainly calculates the distance between each pair of points to assess the similarity between two datasets. For instance, Hausdorff distance [12, 46, 28, 31, 24, 43, 59] measures the similarity between two datasets by calculating the maximum distance from any point in one dataset to its nearest point in the other dataset. Specifically, Nutanong et al. [46] derived the lower and upper bounds of Hausdorff and combined them with the branch-and-bound strategy to improve efficiency. Degbelo et al. [28] conducted experiments to verify that the Hausdorff distance has higher precision than the overlapped area in dataset search. However, the complexity of the existing upper and lower bounds is still high, and we need more efficient bound estimation to speed up the Hausdorff computation. In addition, earth mover’s distance [70, 41, 42] is a more fine-grained point set similarity measure compared to Hausdorff. However, it is a more expensive distance as it has a cubic time complexity.
II-B Spatial Data Point Search
@@
Range-based search.  As one of the fundamental operators in spatial databases, the goal of range-based search is to find all objects falling within the query region. For example, Hariharan et al. [36] constructs a location-based search engine, which supports users to specify a query range, and then all points that fall into the query region are returned to the users. Zacharatou et al. [78] proposed a spatial bundled dataset search over different domains using spatial ranges. In the spatial dataset search system Auctus [19, 1], range queries are supported by returning spatial datasets that overlap with the user-specified query region. Papadopoulos et al. [50] also studied multiple-range query optimization strategies to improve search efficiency by grouping multiple queries.
@@
$k$ Nearest Neighbor search.  The $k$NN [75, 22, 53, 33, 47] is defined as finding $k$ points in a given set that are closest to a given query point. The $k$NN queries have been extensively studied in the literature. To find the $k$ closest neighbor for any given query point, we can utilize the fundamental $k$NN search approach, which is often called the brute force (BF) method. For the given query points, it scans the entire dataset to find the $k$ closest points based on the distances between the query point and all other data points, which is computationally intensive. Thus, to efficiently accelerate the $k$NN search, a variety of indexes with different heuristic improvements have been proposed, such as R tree [34], K-Dimensional tree (KD-tree) [58], and Ball-tree [48]. While these indexes can be implemented to find $k$ nearest neighbor points for a given point, they cannot support both spatial dataset search and spatial data point search, which is the focus of this work.
@@ III-A Data Model
For ease of presentation, we use 2-dimensional points for illustration in the subsequent descriptions. However, all search operations we support can be easily extended to multi-dimensional datasets [51], as shown in Section VII. .
@@
$\small D.b^{\uparrow}[i]=\max_{p\in D}p[i],\quad D.b^{\downarrow}[i]=\min_{p\in D}p[i],wrt.,i\in\{1,2\}$ (1)
@@
When measuring the similarity between two datasets, existing systems [28, 27] mainly use the intersecting area between the MBR of datasets to measure the overlap ratio. To enable a more fine-grained measure of overlap, we introduce a structure called z-order signature to represent each spatial dataset. Specifically, given an integer $\theta$, we partition the entire space containing the spatial data repository into a grid $\mathcal{C}_{\theta}$ of $2^{\theta}\times 2^{\theta}$ equally-sized cells [77, 18, 54], where $\theta$ is referred to as the resolution.
Each cell, denoted as $c$, can be treated as an element. The coordinates of cell $c$ can be converted into a unique non-negative integer $c.id$ by interleaving the binary representations of its coordinate values [52, 78, 57, 70]. This transformation yields consecutive IDs in the range $[0,2^{\theta}\times 2^{\theta}-1]$. In this way, a spatial dataset $D$ can be represented as a sorted integer set consisting of a sequence of cell IDs. The formal definition of z-order signature is as follows.
@@ III-B Similarity Measures
III-B1 Overlap-based Similarity Measures
@@
(Intersecting Area (IA)) Given a query dataset $Q$ and a spatial dataset $D$, the IA between $Q$ and $D$ is the size of the intersecting area of two given datasets’ MBR, i.e., IA$(Q,D)=\prod_{i=1}^{2}l(i,Q,D),$ where $l(i,Q,D)$ denotes the intersecting length of two MBRs in $i$-th dimension [84].
@@
III-B2 Point-wise Similarity Measure
From the above definitions, we can observe that these two similarity measures only work for overlapping datasets. To address this, we introduce a more fine-grained measure, Hausdorff distance [46], which considers the distance between each pair of points for two datasets.

### v3

@@ Abstract.
copyright: acmcopyrightdoi: 10.475copyright: acmcopyrightjournalyear: 2026doi: XXXXXXX.XXXXXXX
PVLDB Reference Format:
A Unified Approach for Multi- Granularity Search over Spatial Datasets. PVLDB, 19(X): XXX-XXX, 2026.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For™ any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. X ISSN 2150-8097.
doi:XX.XX/XXX.XX
1. Introduction
The development of sensor technology, GPS-enabled mobile devices, and wireless communication brings prosperity to spatial data, and the emergence of a large number of spatial data has fueled the demand for searching spatial data (Degbelo2019, ; Bouros2019, ; Nandan2014, ; Zacharatou2019, ; Wang2019d, ; YangWZ2022, ; DeFernandesVasconcelos2017, ). Since a spatial dataset usually consists of a set of spatial data points (YangWZ2022, ), various spatial dataset search (YangWZ2022, ; Li0SWYY24, ; LiPY24, ) and data point search systems (foursquare, ; ChenCCT15, ; Dong2019b, ) have been developed. For example, Foursquare (foursquare, ) helps users find the best location points in any city around the world. OpenStreetMap (openStreetMap, ) provides search functionality over a vast collection of spatial datasets contributed by users worldwide. However, these systems primarily support keyword-based searches limited to published dataset metadata (Chirigati2021, ; ZhongLZ23, ; ChenLS2020, ).
The published spatial metadata is often missing or incomplete (auctus, ), rendering keyword-based queries insufficient for helping users locate datasets or data points. Therefore, various query operations that rely on raw data have been developed to identify spatial datasets and data points. For example, existing spatial dataset search engines treat a dataset as the smallest return unit. In contrast, in the existing spatial data point search engines, a data point is the smallest return unit. In what follows, we use a traffic analysis example to introduce these two types of searches.
Figure 1. The illustration of our multi-granularity search research over spatial datasets, which seamlessly integrates the dataset search (including (1) and (2) based on three distance metrics) and data point search (including (3) and (4)).
@@ 1. Introduction
Consider a scenario where traffic analysts want to analyze traffic patterns in a metropolitan area. Firstly, they can define a query range of $R$ (auctus, ; Zacharatou2019, ; BrinkhoffKS93, ) to locate all relevant traffic datasets within that area for analysis, as illustrated on the left side of Figure1(a). Then, to identify the most efficient commuting routes, the analysts can input a specific traffic dataset $Q$ and search for datasets similar to $Q$ based on defined similarity measures (YangWZ2022, ; Wang2017, ; Wangsheng2018, ) for data analysis, illustrated on the right side of Figure 1(a). This search manner is known as exemplar search (Rezig2021, ; YewLML23, ; WangMGBCDSW24, ; Li0SWYY24, ; Mottin2014a, ; YangZ0H0020, ). Furthermore, if the analysts want to find points of interest, such as gas stations or rest areas along the searched routes $D$, they can specify a query range $R$ along with the relevant traffic dataset to retrieve these points, as depicted on the left side of Fig.1(b)(foursquare, ; Wang2019d, ; ZHANGSongnian2021, ; Qi2020, ). Additionally, on the right side of Figure1(b), the analysts can identify the nearest gas stations for each location in the dataset $Q$ (Qi2020, ; ZHANGSongnian2021, ; Dong2019b, ; ChoudhuryCSC16, ; CorralMTV2000, ), which helps commuters find convenient stops along their routes (YinWWCZ17, ; yao2016poi, ).
@@
What is the Problem? Despite the efforts made in the field of spatial data search systems, existing systems focus mainly on one type of search. There is currently no system that has successfully integrated all types of searches into a unified search framework. In addition, the existing systems, such as Auctus (auctus, ) and OpenStreetMap (openStreetMap, ), are relatively simple in spatial data search and do not take into account the similarity measurement of spatial location distribution characteristics. Our study, depicted in Figure 1, considers all searches at both the coarse-grained spatial dataset level and the more fine-grained data point level. In spatial dataset search, we integrate range-based dataset search and top-$k$ exemplar dataset search based on several distance metrics.
@@
Figure 2. The illustration of three distance metrics, where (a) shows that the IA of $Q$ and $D_{1}$ is the size of their minimum bounding rectangles’ overlapping areas, i.e., the rectangle filled with red lines, (b) shows that the GBO of $Q$ and $D_{1}$ is 1 since there is one overlapping cell between $Q$ and $D_{1}$, and (c) shows that the Haus of $Q$ and $D_{1}$ is the maximum nearest neighbor distance, i.e., the distance from the point $p_{1}$ of $Q$ to the point $p_{1}^{\prime}$ of $D_{1}$.
@@
As shown in Figure 2, the intersecting area (IA(Larson2011, ; Degbelo2019, ) is useful when assessing the overlap between public transportation routes, which can help identify congestion points to optimize traffic management. Compared to IA, grid-based overlap (GBO(Nandan2014, ) can capture the density of traffic in the urban grid in a finer granularity and optimize traffic routes. In addition, calculating the Hausdorff distance (Haus(Taha2015, ; Nutanong2011, ; Frontiera2008, ) can help analysts understand the accessibility of bus stops for residents (the formal definition is in Section 3.2). Moreover, building upon the outcomes of the dataset search, we further enhance our capabilities to facilitate a finer-grained data point search process, wherein the results obtained from the dataset search are fed to the subsequent data point search.
@@
What are the Challenges? Our goal is to design a multi-granularity spatial data search system that seamlessly supports coarse-grained dataset search based on multiple distance metrics and fine-grained data point search. There are several challenges: Firstly, A naïve way is to build separate indexes for each query type. However, this method incurs significant costs in both storage and index construction time. Furthermore, the presence of outliers in spatial datasets, such as erroneous data points caused by GPS errors, can severely affect the accuracy and efficiency of similarity computation (Wang2012, ). Finally, although the index can quickly locate data matching query criteria, it does not effectively reduce the number of candidate points to examine. Thus, there is an urgent need for efficient acceleration strategies to speed up a rich set of searches.
@@
How the Challenges are Addressed? To address the challenges above, we develop a multi-granularity spatial data search system, called Spadas. Specifically, we design a unified index (Wang2018a, ; Hoang-Vu2016, ) that can effectively organize both datasets and data points, enabling efficient pruning for various search scenarios. In addition, to further enhance system performance, we propose a set of novel and highly efficient pruning mechanisms, such as efficient bound estimation or error-bounded approximation techniques, ensuring that irrelevant candidates are filtered early in the process, thereby reducing computational overhead and improving overall system performance. Our main contributions are summarized as follows.
@@
  • We propose a multi-granularity spatial data search system, called Spadas, to efficiently support coarse-grained dataset search based on multiple distance metrics and fine-grained data point search (see Section 4). We also implement an online Spadas system [^1] http://swangdbs.com/prototype/spadas/ that can be accessed by users.
@@
  • To support both searches with different query types and distance metrics, we specifically design a unified index to enhance the efficiency of data search operations by efficiently organizing spatial data, while mitigating the influence of outliers on the accuracy of similarity measures between datasets (see Section 5).
@@
  • To accelerate multiple types of data search operations, we propose a range of pruning mechanisms over the unified index, including tighter lower and upper bounds, approximation technique with error bound, and batch pruning, that effectively filter candidates and thus significantly reduce computational overhead (see Section 6).
@@
  • We conduct extensive experiments on six real-world data repositories, and the results show that our proposed Spadas can achieve orders of magnitude speedup than the state-of-the-art algorithms in dataset search and data point search. The case studies validate the effectiveness of our search system (see Section 7).
2. Related Work
@@
2.1. Spatial Dataset Search
@@
Overlap-based Similarity Search.  For a user with a query dataset, the overlap-based similarity search ranks candidate datasets based on the size of their overlap area. Computing the overlapping area between two Minimum Bounding Rectangles (MBRs) is a commonly used and straightforward measure. Frontiera et al. (Frontiera2008, ) proposed several variants by considering the query and the dataset’s range. Vasconcelos et al. (DeFernandesVasconcelos2017, ) and Degbelo et al. (Degbelo2019, ) applied it as the main ranking metric for an open government dataset search portal. More recently, with the rise of grid partition technology (Zacharatou2019, ; Winter2019, ), a more fine-grained similarity measure based on the number of overlapping grid cells began to appear, i.e., modeling each dataset as a grid with points inside by dividing the space equally and using the intersection of the grids as the similarity.
@@
Point Set Similarity Search.  Point set similarity search mainly calculates the distance between each pair of points to assess the similarity between two datasets. For instance, Hausdorff distance (Adelfio2011, ; Nutanong2011, ; Degbelo2019, ; Gao2014, ; Chen2017g, ; Li2017e, ; Taha2015, ) measures the similarity between two datasets by calculating the maximum distance from any point in one dataset to its nearest point in the other dataset. Specifically, Nutanong et al. (Nutanong2011, ) derived the lower and upper bounds of Hausdorff and combined them with the branch-and-bound strategy to improve efficiency. Degbelo et al. (Degbelo2019, ) conducted experiments to verify that the Hausdorff distance has higher precision than the overlapped area in dataset search. However, the complexity of the existing upper and lower bounds is still high, and we need more efficient bound estimation to speed up the Hausdorff computation. In addition, earth mover’s distance (YangWZ2022, ; Li0SWYY24, ; LiPY24, ) is a more fine-grained point set similarity measure compared to Hausdorff. However, it is a more expensive distance as it has a cubic time complexity.
2.2. Spatial Data Point Search
@@
Range-based search.  As one of the fundamental operators in spatial databases, the goal of range-based search is to find all objects falling within the query region. For example, Hariharan et al. (HariharanHLM07, ) constructs a location-based search engine, which supports users to specify a query range, and then all points that fall into the query region are returned to the users. Zacharatou et al. (Zacharatou2019, ) proposed a spatial bundled dataset search over different domains using spatial ranges. In the spatial dataset search system Auctus (Chirigati2021, ; auctus, ), range queries are supported by returning spatial datasets that overlap with the user-specified query region. Papadopoulos et al. (PapadopoulosM98, ) also studied multiple-range query optimization strategies to improve search efficiency by grouping multiple queries.
@@
$k$ Nearest Neighbor search.  The $k$NN (YuCWS07, ; ChenCCT15, ; Qi2020, ; G2010, ; NutanongS13, ) is defined as finding $k$ points in a given set that are closest to a given query point. The $k$NN queries have been extensively studied in the literature. To find the $k$ closest neighbor for any given query point, we can utilize the fundamental $k$NN search approach, which is often called the brute force (BF) method. For the given query points, it scans the entire dataset to find the $k$ closest points based on the distances between the query point and all other data points, which is computationally intensive. Thus, to efficiently accelerate the $k$NN search, a variety of indexes with different heuristic improvements have been proposed, such as R tree (Guttman84, ), K-Dimensional tree (KD-tree) (Sproull91, ), and Ball-tree (Omohundro1989, ). While these indexes can be implemented to find $k$ nearest neighbor points for a given point, they cannot support both spatial dataset search and spatial data point search, which is the focus of this work.
@@ 3.1. Data Model
For ease of presentation, we use 2-dimensional points for illustration in the subsequent descriptions. However, all search operations we support can be easily extended to multi-dimensional datasets (patro2015normalization, ), as shown in Section 7. .
@@
(1) $\small D.b^{\uparrow}[i]=\max_{p\in D}p[i],\quad D.b^{\downarrow}[i]=\min_{p\in D}p[i],wrt.,i\in\{1,2\}$
@@
When measuring the similarity between two datasets, existing systems (Degbelo2019, ; DeFernandesVasconcelos2017, ) mainly use the intersecting area between the MBR of datasets to measure the overlap ratio. To enable a more fine-grained measure of overlap, we introduce a structure called z-order signature to represent each spatial dataset. Specifically, given an integer $\theta$, we partition the entire space containing the spatial data repository into a grid $\mathcal{C}_{\theta}$ of $2^{\theta}\times 2^{\theta}$ equally-sized cells (Zacharatou2017, ; Cao2021a, ; QiaoHZWGW20, ), where $\theta$ is referred to as the resolution.
Each cell, denoted as $c$, can be treated as an element. The coordinates of cell $c$ can be converted into a unique non-negative integer $c.id$ by interleaving the binary representations of its coordinate values (Peng2016, ; Zacharatou2019, ; SieranojaF18, ; YangWZ2022, ). This transformation yields consecutive IDs in the range $[0,2^{\theta}\times 2^{\theta}-1]$. In this way, a spatial dataset $D$ can be represented as a sorted integer set consisting of a sequence of cell IDs. The formal definition of z-order signature is as follows.
@@ 3.2. Similarity Measures
3.2.1. Overlap-based Similarity Measures
@@
(Intersecting Area (IA)) Given a query dataset $Q$ and a spatial dataset $D$, the IA between $Q$ and $D$ is the size of the intersecting area of two given datasets’ MBR, i.e., IA$(Q,D)=\prod_{i=1}^{2}l(i,Q,D),$ where $l(i,Q,D)$ denotes the intersecting length of two MBRs in $i$-th dimension (Zomorodian2000, ).
@@
3.2.2. Point-wise Similarity Measure
From the above definitions, we can observe that these two similarity measures only work for overlapping datasets. To address this, we introduce a more fine-grained measure, Hausdorff distance (Nutanong2011, ), which considers the distance between each pair of points for two datasets.